# ZeroPur: Succinct Training-Free Adversarial Purification

### v1

@@ 4.1 Quantitative Evaluation and Comparison
For adversarial defense, adversarial training (AT) involves adversarial examples in the classifier training, and auxiliary-based purification (ABP) introduces an auxiliary function to cooperate with the victim classifier. These methods do not rely on external generative models but still necessitate training auxiliary functions or retraining classifiers. External model-based purification (EBP) fine-tunes or retrains an external generative model to remove adversarial perturbations in adversarial images. Since the performance of EBP is ensured by the modeling capability of generative models, it typically outperforms ABP. In this section, for a fair comparison, we will first compare our method with both AT and ABP and then compare it with EBP. ’Require Training’ in Table 1,  3 and  3 denotes that the methods require retraining external generative models or auxiliary functions or victim classifiers.
Comparison with AT &\&& ABP. Table 1 reports the standard and robust accuracy against AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$). ZeroPur achieves an approximate $10\%$ increase in robust accuracy for ResNet-18, and this improvement surpasses the performance of most other methods using WideResNet-28-10. We also report the comparison on ImageNet1-K in Table 3, where ZeroPur improves robust accuracy by $30\%$ compared to AT and ABP methods. These results also demonstrate that AP can significantly move adversarial images toward the natural image manifold. It is noteworthy to mention that our method consistently presents better performance in small architectures such as ResNet-18 compared to large architectures. This stems from the inclination of our method to ‘fine-tune’ adversarial images, enabling them to approach natural images, whereas larger models often introduce more hindrances to this process.
Comparison with EBP. Training an external generative model to model a transformation from adversarial images to natural images can contribute to an outstanding purification method, but this process is exceedingly computation-consuming. Table 3 and Table 3 report the comparison result between ZeroPur and EBP. The robust accuracy achieved by ZeroPur is only slightly lower than DISCO (Ho et al., 2022) which involves training an external model on ImageNet. Therefore, adjusting ZeroPur a little bit (see the first part of Subsection 4.2) can achieve comparable performance with these EBP methods. Other detailed results are reported in Appendix A.4.
Table 2: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=8/255)$ on CIFAR-10 in comparison with EBP methods.
@@
Table 3: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=4/255)$ on ImageNet-1K.
@@ 4.2 Discussion on ZeroPur
Evaluating ZeroPur on protecting the victim classifier trained with different data augmentation. To further discuss the performance of ZeroPur, we consider three data augmentation strategies (’Vanilla’, ‘Base’, ‘Strong’) to train victim classifiers (See Appendix A.2 for details). ‘Base’ is the most common data augmentation in training classifiers. In the experiment, ZeroPur using a median filter ($3\times 3$ window size) is applied in ’Vanilla’ classifiers and Gaussian blur ($\sigma=1.2$) in ‘Base’ and ’Strong’ classifiers. Table 5 reports the standard and robust accuracy of ZeroPur. We can see that ’Strong’ data augmentation benefits our method. The reason is that such augmentations enable the blurred counterparts to guide adversarial images toward the natural image manifold more effectively, which contributes to precise adaptive projection. The performance of ZeroPur in Table 3 also demonstrates this phenomenon.
[…]
Accuracy ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
CIFAR-10
Standard 82.04 92.56 90.05 91.22 91.81 85.02
Robust 59.38 69.62 83.84 59.16 68.60 82.76
CIFAR-100
Standard 55.45 52.45 67.11 64.89 64.88 66.39
Robust 30.42 41.42 50.66 30.81 34.58 50.25
Table 5: The ablation study on CIFAR-10 using perceptual regularization.
Impact of perceptual regularization. The important of incorporating the perceptual regularization in (Eq.(15)) is revealed by Fig. 5. We can see the robust accuracy of AP experiences a slight decrease in (a) and (b), while the standard accuracy shows significant improvement in (c) and (d). This phenomenon is observed in both ResNet-18 and WideResNet-28-10, indicating the effectiveness of perceptual regularization in AP.
@@
ZeroPur defense against adversarial attacks robust to blurring. ZeroPur utilizes the blurring operator to return adversarial images to the natural image manifold. However, certain attacks, such as DI2-FGSM [49], inherently withstand blurring operations. Indeed, they cannot be entirely immune to ZeroPur. As shown in Table 7, replacing the blurring operator with TVM [13] leads to improved purification. The robust accuracy against AutoAttack increases to $75.96\%$ without training, which outperforms the optimal performance of GS(Blur)+AP by $6.34\%$. This result also demonstrates that our method can enhance any operator capable of destroying adversarial properties, not just blurring.
ZeroPur defense against attacks that bypass the purification module. The efficiency of purification methods can be challenged by the BPDA attack that approximates gradients, bypassing the purification module. Table 7 reports the performance of ZeroPur defense against BPDA attack, where the victim classifiers are WideResNet-28-10 (WRN) and ResNet-18 (R18). We can see that the robust accuracy of ZeroPur decreases, but it remains stable, which suggests that ZeroPur still offers acceptable defense. Since the natural accuracy of ’Vanilla’ WRN on natural images is $71.50\%$, while that of ’Vanilla’ R18 is $57.01\%$, which may result in an approximately $20\%$ gap in evaluation on BPDA (See Table 8 in Appendix A.2).
[…]
Method DI2-FGSM AutoAttack
Vanilla Base Strong Vanilla Base Strong
GS (Blur) 35.77 40.96 66.66 54.37 58.82 79.36
GS (TVM) 38.08 46.79 45.93 56.25 67.81 68.48
GS (Blur) + AP 35.34 59.81 84.41 59.38 69.62 83.84
GS (TVM) + AP 50.51 65.98 66.74 61.75 75.96 78.98
@@ A.2 Implementation Details of Classifiers
Our experiments on CIFAR-10 and CIFAR-100 included ResNet-18 and WideResNet-28-10 trained on three different data augmentation strategies ‘Vanilla’, ‘Base’, and ‘Strong’. Table 8 shows the training settings and natural accuracy. ‘Base’ is the most common data augmentation used by adversarial training methods and adversarial purification methods for better natural accuracy.
@@
Strong 91.08 68.52
WRN-28-10 Vanilla 91.34 71.50
Base 93.83 74.95
Strong 91.09 67.52
@@ A.3 Implementation Details of Adversarial Attacks
AutoAttack [6].  We use AutoAttack to compare with the start-of-the-art methods. The robust classifier for adversarial training methods provided by RobustBench [7] benchmark available at https://robustbench.github.io. The code for adversarial purification methods is provided by their respective papers.
@@ A.4 More Experimental Results
In Table 9 and Table 10, we compare our method with AT, ABP, and EBP methods on CIFAR-10 against $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ threat model, detailing training requirements for each method. We also report the performance of ZeroPur when victim classifiers use strong data augmentation (ZeroPur). Comparing all methods, ZeroPur always achieves the second-best performance. However, when comparing ABP methods, ZeroPur achieves optimal performance.
Table 9: Standard and robust accuracy (%) against AutoAttack $\ell_{\infty}(\epsilon=8/255)$ on CIFAR-10 in comparison with three types methods, obtained by different classifier architectures.
Defense Type Method Training Architecture Standard Acc Robust Acc
@@
(Addepalli et al., 2022) Classifier ResNet-18 85.71 82.90
@@
(Wang et al., 2023) Classifier WRN-28-10 92.44 67.31
@@
(Hwang et al., 2023) Auxiliary branch WRN-34-10 87.02 56.63
@@
Table 10: Standard and robust accuracy (%) against AutoAttack $\ell_{2}(\epsilon=0.5)$ on CIFAR-10 in comparison with three types methods, obtained by different classifier architectures.
Defense Type Method Training Architecture Standard Acc Robust Acc
@@
[…]
Operator Level ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
Median $3\times 3$ 59.38 72.27 60.74 59.16 67.56 70.01
$5\times 5$ 34.17 64.81 44.63 39.81 58.41 50.49
$7\times 7$ 27.25 54.41 34.68 28.94 48.26 40.58
Gaussian $\sigma=0.6$ 38.82 65.43 64.92 53.21 58.31 73.79
$\sigma=1.2$ 26.20 69.62 83.84 27.97 68.60 82.76
$\sigma=1.8$ 20.45 60.67 67.89 26.82 59.91 78.82
[…]
Operator Level ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
Median $3\times 3$ 82.04 92.34 90.90 91.22 91.81 90.65
$5\times 5$ 83.73 91.87 90.92 91.06 91.34 88.97
$7\times 7$ 83.66 91.56 90.87 90.97 90.34 87.70
Gaussian $\sigma=0.6$ 83.57 91.64 90.72 91.09 90.02 90.68
$\sigma=1.2$ 83.63 92.56 90.05 91.14 91.81 85.02
$\sigma=1.8$ 83.69 92.36 90.79 51.26 35.95 80.27
Table 11 and Table 12 report the robust accuracy and standard accuracy of ZeroPur using different blur operators in GS. We use 5 different blurring operators including median filters with windows size $3\times 3$, $5\times 5$, and $7\times 7$ and Gaussian blurring kernel with $\sigma=0.6$, $1.2$, and $1.8$. The results show that if victim classifiers are not trained on strong data augmentation, low-quality embeddings of images from excessive blurring will hinder ZeroPur’s ability to move adversarial images. On the contrary, the victim classifiers trained on strong data augmentation can recognize these embeddings, enabling ZeroPur to provide precise movement direction.
We provide visual examples in Fig. 5 to illustrate several results and the visualization of this process, where diff_$I_{a}$_$I_{b}$ = $\mathrm{abs}(I_{a}-I_{b})$ represents the difference between the two images. The resemblance between diff_gs_nat and diff_adv_nat outlines suggests that Guided Shift can effectively reverse the adversarial perturbation at a coarse level. Furthermore, the close correspondence in detail between diff_adv_nat and diff_ap_nat demonstrates the precise approximation of this perturbation by Adaptive Projection. diff_ap_adv quantifies the distance of the adversarial image moved by ZeroPur.
(a) Examples purified by GS
(b) Examples purified by GS + AP
Figure 5: Visual examples of ZeroPur against $\ell_{\infty}$ threat model ($\epsilon=4/255$) on ImageNet-1K. The central two columns depict the results of GS and AP, while the red label denotes error prediction and the green label denotes correct prediction. diff_$I_{a}$_$I_{b}$ = $\mathrm{abs}(I_{a}-I_{b})$ represents the difference between the two images.

### v2

@@ 4.1 Quantitative Evaluation and Comparison
For adversarial defense, adversarial training (AT) involves adversarial examples in the classifier training, and auxiliary-based purification (ABP) introduces an auxiliary function to cooperate with the victim classifier. These methods do not rely on external generative models but still necessitate training auxiliary functions or retraining classifiers. External model-based purification (EBP) fine-tunes or retrains an external generative model to remove adversarial perturbations in adversarial images. Since the performance of EBP is ensured by the modeling capability of generative models, it typically outperforms ABP. In this section, for a fair comparison, we will first compare our method with both AT and ABP and then compare it with EBP. In Table 1,  3, and  3, the terms ’Ext’, ’Aux’, and ’Cls’ under ’Require Training’ indicate whether the methods require retraining external generative models, auxiliary functions, or victim classifiers.
Comparison with AT &\&& ABP. Table 1 reports the standard and robust accuracy against AutoAttack $\ell_{\infty}$ ($\epsilon=8/255$). ZeroPur achieves consistent and significant improvements in robust accuracy, particularly on ResNet-18, with robust accuracies of $69.62\%$ and $41.42\%$ on CIFAR-10 and CIFAR-100, respectively. Additionally, we report the comparison on ImageNet-1K in Table 3, where ZeroPur improves robust accuracy by $30.06\%$ compared to AT and ABP methods, representing a substantial gain. These results demonstrate that ZeroPur effectively guides adversarial images toward the natural image manifold. Notably, our method performs better on smaller architectures like ResNet-18, likely because larger models have more complex feature spaces, making them more sensitive to attacks, which increases the difficulty of mapping adversarial images to natural ones using our method.
Comparison with EBP. Although generative model-based purification methods achieve outstanding performance due to their large scale and high computation, our method still achieves state-of-the-art or comparable performance. Table 3 and Table 3 report the comparison result between ZeroPur and EBP. For CIFAR-10, ZeroPur achieves robust accuracy comparable to existing EBP defense methods, and a slightly tweaked ZeroPur (see the first part of Section 4.2) can achieve state-of-the-art performance over EBP methods. On ImageNet, our method achieves SOTA robustness, but with a slight decrease in natural accuracy. For example, even compared to the latest Mimicdiffusion [43], our method demonstrates a $5.26\%$ improvement in robust accuracy with only a $3.15\%$ loss in clean accuracy. Figure 6 shows the visualizations and differences resulting from applying the two-step process of ZeroPur to adversarial examples. The reduced difference between the purified and natural images, along with the increased difference from adversarial images, demonstrates the effectiveness of our method. More experimental results and visualizations (including failure cases) are provided in Appendix A.4.
Table 2: Comparison with EBP methods on CIFAR-10 under AutoAttack $\ell_{\infty}(\epsilon=8/255)$.
@@
Table 3: Comparison on ImageNet-1K under AutoAttack $\ell_{\infty}(\epsilon=4/255)$.
@@ 4.2 Discussion on ZeroPur
Evaluating ZeroPur on protecting the victim classifier trained with different data augmentation. To further discuss the performance of ZeroPur, we consider three data augmentation strategies (’Vanilla’, ‘Base’, ‘Strong’) to train victim classifiers (See Appendix A.2 for details). ‘Base’ is the most common data augmentation in training classifiers. In the experiment, ZeroPur using a Median filter ($3\times 3$ window size) is applied in ’Vanilla’ classifiers and Gaussian blur ($\sigma=1.2$) in ‘Base’ and ’Strong’ classifiers, please refer to Table 6 in Appendix A.4. Figure 5(a) and Figure 5(b) report the standard and robust accuracy of ZeroPur under three different training recipes for ResNet-18 and WideResNet-28-10, respectively. It can be observed that the ’Strong’ data augmentation benefits our method. The reason for this is that models trained with such augmentation exhibit better perceptual capability for blurred images, enabling them to more effectively guide adversarial images towards the natural image manifold and facilitate precise adaptive projection. The performance of ZeroPur in Table 3 also demonstrates this phenomenon.
(a) ResNet-18
(b) WideResNet-28-10
(c) Ablation of PR
(d) Blur vs. TVM
Figure 5: SA and RA (%) of classifiers trained with three-level data augmentation on (a) ResNet-18 and (b) WideResNet-28-10. (c) Ablation Study of perceptual regularization on CIFAR-10. (d) Robust accuracy against DI2-FGSM and AutoAttack on CIFAR-10 by Blur and TVM.
Impact of perceptual regularization. Figure 5(c) illustrates the impact of perceptual regularization in Eq.(15) on standard accuracy. The comparison between models with and without PR (red vs. green) reveals that incorporating PR leads to a notable improvement in standard accuracy with almost no loss in robust accuracy. This trend is consistent across different architectures, ResNet-18 (R18) and WideResNet-28-10 (WRN), confirming the effectiveness of PR in the adaptive projection stage.
Figure 6: Visual examples of ZeroPur against $\ell_{\infty}$ threat model ($\epsilon=4/255$) on ImageNet-1K. The red denotes error prediction and the green denotes correct prediction. $\Delta_{\mathrm{A}\text{-}\mathrm{B}}$ represents the normalized absolute difference between A and B.
@@
ZeroPur defense against adversarial attacks robust to blurring. ZeroPur employs blurring to project adversarial images back onto the natural image manifold. Although certain attacks, such as DI2-FGSM[54], exhibit inherent resistance to blurring, they are still not entirely immune to ZeroPur. As shown in Figure 5(d), replacing the blurring operator with TVM [12] leads to improved purification. The robust accuracy against AutoAttack increases to $75.96\%$ without training, which outperforms the optimal performance of GS(Blur)+AP by $6.34\%$. This result also demonstrates that ZeroPur is not limited to blurring and can generalize to other operators.
@@
ZeroPur defense against attacks that bypass the purification module. The efficiency of purification methods can be challenged by the BPDA attack that approximates gradients, bypassing the purification module. Table 4.2 reports the performance of ZeroPur defense against BPDA attack. We can observe that the robust accuracy of ZeroPur decreases, but remains stable, indicating that ZeroPur provides acceptable defense. As mentioned earlier, the larger model WRN consistently exhibits slightly worse robustness against BPDA compared to the smaller model R18, while achieving better natural accuracy than R18 (See Table 5 in Appendix A.2).
@@ A.2 Implementation Details of Classifiers
Our experiments on CIFAR-10 and CIFAR-100 included ResNet-18 and WideResNet-28-10 trained on three different data augmentation strategies: ‘Vanilla’, ‘Base’, and ‘Strong’. Table 5 shows the training settings and natural accuracy. ‘Base’ is the most common data augmentation used by adversarial training methods and adversarial purification methods for better natural accuracy.
@@
Strong 91.08 68.53
WRN-28-10 Vanilla 91.34 71.57
Base 93.83 74.59
Strong 91.10 67.65
@@ A.3 Implementation Details of Adversarial Attacks
AutoAttack [4].  We use AutoAttack to compare with the start-of-the-art methods. The robust classifier for adversarial training methods provided by RobustBench [5] benchmark is available at https://robustbench.github.io. The code for adversarial purification methods is provided by their respective papers.
@@
Table 6: SA / RA (%) on CIFAR-10 with different augmentations under varying blurring operators. The bold indicates the best harmonic mean of SA and RA, measured by the F1 Score.
Operator Level ResNet-18 WRN-28-10
Vanilla Base Strong Vanilla Base Strong
Median $3\times 3$ 82.04/ 59.38 92.34/ 72.27 90.90/ 60.74 91.22/ 59.16 91.81/ 67.56 90.65/ 70.01
$5\times 5$ 83.73/ 34.17 91.87/ 64.81 90.92/ 44.63 91.06/ 39.81 91.34/ 58.41 88.97/ 50.49
$7\times 7$ 83.66/ 27.25 91.56/ 54.41 90.87/ 34.68 90.97/ 28.94 90.34/ 48.26 87.70/ 40.58
Gaussian $\sigma=0.6$ 83.57/ 38.82 91.64/ 65.43 90.72/ 64.92 91.09/ 53.21 90.02/ 58.31 90.68/ 73.79
$\sigma=1.2$ 83.63/ 26.20 92.56/ 69.62 90.05/ 83.84 91.14/ 27.97 91.81/ 68.80 85.02/ 82.76
$\sigma=1.8$ 83.69/ 20.45 92.36/ 60.67 90.79/ 67.89 51.26/ 26.82 35.95/ 59.91 80.27/ 78.82
@@ A.4 More Experimental Results
In Table 7 and Table 8, we compare our method with AT, ABP, and EBP methods on CIFAR-10 against $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$ threat model, detailing training requirements for each method. We also report the performance of ZeroPur when victim classifiers use strong data augmentation (ZeroPur). Among all methods, ZeroPur generally achieves comparable standard and robust accuracy to any other method, while its enhanced version ZeroPur consistently attains the best robust accuracy.
To evaluate the effect of training strategies and blur types on ZeroPur, Table 6 presents its standard and robust accuracy under different augmentations (Vanilla, Base, Strong) using Median filters ($3\times 3$, $5\times 5$, $7\times 7$) and Gaussian blurs ($\sigma=0.6$, $1.2$, $1.8$). The results show that if the victim classifiers are not trained with strong data augmentation, the low-quality embeddings induced by stronger blur operators will hinder ZeroPur’s ability to guide adversarial images toward the natural manifold. In contrast, victim classifiers trained with strong data augmentation inherently possess the ability to recognize blurred images, enabling ZeroPur to benefit from high-intensity blurring. For commonly used base classifiers, they generally perform well under moderate blur intensity.
We provide additional visualization examples in Figure 7 to further validate the effectiveness and interpretability of ZeroPur. The progression from adversarial images to GS-purified images and then to AP-purified images shows decreasing differences from natural images (gradually transitioning from yellow to blue), indicating the effectiveness of our method in purification. We observe that in cases where GS successfully purifies adversarial samples, the difference between GS-purified and natural images (denoted as $\Delta_{\text{GS-Nat.}}$) displays scattered, non-semantic noise. In contrast, for cases where GS fails to purify, the difference $\Delta_{\text{GS-Nat.}}$ still shows semantic perturbations. Furthermore, the similarity between $\Delta_{\text{AP-Adv.}}$ and $\Delta_{\text{Adv.-Nat.}}$ suggests that our method (GS+AP, i.e., ZeroPur) focuses purification mainly on regions influenced by adversarial perturbations. However, there are still some failure cases. It is not difficult to observe that the inconsistency between the predictions of Adv. and GS already appears at the beginning. This highlights the importance of the initial purification direction, which may be the key to ZeroPur’s success.
Table 7: Comparison with three types methods on CIFAR-10 under AutoAttack $\ell_{\infty}(\epsilon=8/255)$.
Defense Type Method Training Architecture SA(%) RA(%)
@@
(Addepalli et al., 2022) Classifier ResNet-18 85.71 52.90
@@
(Wang et al., 2023b) Classifier WRN-28-10 92.44 67.31
(Li and Spratling, 2023) Classifier ResNet-18 83.45 52.52
(Cui et al., 2024) Classifier WRN-28-10 92.16 67.73
@@
(Hwang et al., 2023) Auxiliary branch WRN-34-10 82.22 56.63
@@
Table 8: Comparison with three types methods on CIFAR-10 under AutoAttack $\ell_{2}(\epsilon=0.5)$.
Defense Type Method Training Architecture SA(%) RA(%)
@@
Figure 7: Visual examples of ZeroPur against $\ell_{\infty}$ threat model ($\epsilon=4/255$) on ImageNet-1K. The red denotes error prediction and the green denotes correct prediction. $\Delta_{\mathrm{A}\text{-}\mathrm{B}}$ represents the normalized absolute difference between A and B. (Zoom in for a clearer view.)